OpenStudy (hanna1):

plz help!! Find the value of r so that the slope between the points is 4. (6, -2) and (r, -6) a. -5 b. 5 c. -42 d. 4

OpenStudy (avi13):

B 5 as 4=(-6-(-2))/r-6=-4/r66 therefore r-6=-1 r=5

OpenStudy (amorfide):

\[m=\frac{ y_2{}-y_{1} }{ x_{2}-x_{1} }\] where a coordinate is \[(x_{1},y_{1})\] since you have 2 coordinates, your other will be \[(x_{2},y_{2})\] so use these values and substitute to find m, which is the slope we know m=4 so you have \[4=\frac{ y_2{}-y_{1} }{ x_{2}-x_{1} }\]

OpenStudy (amorfide):

then you would solve it for r